# Resource-Efficient Federated Multimodal Learning via Layer-wise and Progressive Training

### v1

@@ Abstract
Combining different data modalities enables deep neural networks to tackle complex tasks more effectively, making multimodal learning increasingly popular. To harness multimodal data closer to end users, it is essential to integrate multimodal learning with privacy-preserving training approaches such as federated learning (FL). However, compared to conventional unimodal learning, multimodal setting requires dedicated encoders for each modality, resulting in larger and more complex models that demand significant resources. This presents a substantial challenge for FL clients operating with limited computational resources and communication bandwidth. To address these challenges, we introduce LW-FedMML, a layer-wise federated multimodal learning approach, which decomposes the training process into multiple steps. Each step focuses on training only a portion of the model, thereby significantly reducing the memory and computational requirements. Moreover, FL clients only need to exchange the trained model portion with the central server, lowering the resulting communication cost. We conduct extensive experiments across various FL scenarios and multimodal learning setups to validate the effectiveness of our proposed method. The results demonstrate that LW-FedMML can compete with conventional end-to-end federated multimodal learning (FedMML) while significantly reducing the resource burden on FL clients. Specifically, LW-FedMML reduces memory usage by up to 2.7×2.7$\times$2.7 ×, computational operations (FLOPs) by 2.4×2.4$\times$2.4 ×, and total communication cost by 2.3×2.3$\times$2.3 ×. We also introduce a progressive training approach called Prog-FedMML. While it offers lesser resource efficiency than LW-FedMML, Prog-FedMML has the potential to surpass the performance of end-to-end FedMML, making it a viable option for scenarios with fewer resource constraints.
@@ I Introduction
Deep neural networks have become an integral part of many practical applications in our daily lives. While these models have traditionally been trained on a single modality, recent advancements in the field are driving a shift towards multimodal learning [1, 2, 3]. This transition is motivated by the realization that leveraging multiple types of data simultaneously can significantly enhance the performance of the models [4]. Just as animal intelligence relies on multiple senses to perceive their environment, combining different data modalities allows models to tackle complex tasks more effectively. Integrating diverse data sources, such as vision and audio, ensures that more information is available to the model, thereby improving its perception and decision-making capabilities [5].
Multimodal learning based on instance discrimination methods aligns representations of different modalities within the same data sample, thereby learning transferable representations [3, 6]. On the other hand, supervised multimodal learning aims to fuse information from various modalities to enhance the predictive performance of the model [7, 8, 9]. As mobile and IoT devices become increasingly powerful, they are integrated with a broader range of sensors [10]. For instance, smartphones are equipped with cameras, gyroscopes, accelerometers, and more, while smart home IoT devices can include an array of modular sensors for various functions. These edge devices capture and store valuable real-world multimodal data, which is crucial for training AI models for practical applications. However, conventional multimodal learning methods require collecting this data into a central repository for training, raising data privacy concerns.
To harness the distributed data while also considering privacy issues, numerous studies have explored decentralized training approaches like federated learning (FL) [11, 12, 13, 14]. In an FL system, the central server collects model parameters trained on the private data of edge devices (known as FL clients) rather than the data itself [11]. Such a mechanism effectively guarantees the confidentiality of data held by different parties involved in the training process. However, this also means that the FL training process relies heavily on the computation and communication capabilities of FL clients, as they are responsible for training the model [15].
@@
Figure 1: Overview of the federated learning process. (i) The global model $M$ is distributed to clients. (ii) Each client performs training on the local dataset $D^{n}$. (iii) The local model $M^{n}$ is sent back to the server. (iv) The server performs aggregation.
During training, reducing the batch size may appear to be a straightforward method to lower the computational demands on resource-constrained devices. However, for instance discrimination-based multimodal learning approaches (e.g., CLIP [3]), a small batch size may compromise the effectiveness of the contrastive loss, degrading the quality of learned representations [17, 18]. These methods benefit from a larger batch size to increase the number of negative samples, which is essential for the contrastive loss to perform well. Furthermore, simply reducing the batch size only addresses the computational requirements but fails to alleviate the communication burden on FL clients, as they still need to exchange the same model size with the server.
@@
Our contributions are summarized as follows: (i) We introduce LW-FedMML, a layer-wise federated multimodal learning strategy designed to address the resource limitations of FL clients engaged in multimodal tasks. LW-FedMML can match the performance of conventional end-to-end training while significantly reducing resource requirements. (ii) We also explore a progressive training approach, referred to as Prog-FedMML. Although Prog-FedMML does not offer the same level of resource efficiency as LW-FedMML, it has the potential to outperform conventional end-to-end training while maintaining similar memory usage and lowering computational and communication costs. (iii) To validate the effectiveness of our proposed methods, we conducted extensive experiments across various FL scenarios and multimodal learning setups.
(a) Instance discrimination-based multimodal learning.
(b) Supervised multimodal learning.
Figure 2: Common multimodal learning approaches with modalities $a$ and $b$. $M$ denotes the model, and $x$ denotes the input sample. (a) The instance discrimination-based approach focuses on aligning the representations of diverse modalities within a given data sample. Negative samples are omitted in the figure for clarity. (b) The supervised approach aims to fuse information from various modalities, enabling the model to make more informed predictions. Here, $H_{\text{sup}}$ can be any task-specific prediction head.
Figure 3: Overview of LW-FedMML across different stages. In stage $s$, the active layers $L_{a}^{s}$ and $L_{b}^{s}$ are depicted in green within encoders $F_{a}$ and $F_{b}$, where $a$ and $b$ represent different modalities. Prior layers within the encoders are frozen, as indicated in gray.
Figure 4: Overview of LW-FedMML for supervised setting at stage $s$.
@@ II-A Federated Learning
A federated learning (FL) process typically involves a central server and a set of FL client devices [11, 19, 20]. These devices are often found close to end-users, capturing and holding valuable data for model training. Meanwhile, the central server oversees and monitors the overall FL procedure. Given $N$ clients, the goal of FL is to train a global model $M$ by solving the following optimization problem:
@@
with $|D^{n}|$ and $|D|$ representing the number of data samples in $D^{n}$ and $D$, respectively. As illustrated in Fig. 1, four key steps comprise a standard FL communication round. (i) The central server distributes the global model $M$ to each participating client $n\in[1,N]$. (ii) Each client $n$ trains the received model on its local dataset $D^{n}$ to produce the local model $M^{n}$. (iii) The trained local models $[M^{1},\dots,M^{N}]$ are sent back to the server. (iv) The server aggregates the received local models and updates the global model. The FedAvg algorithm is commonly used for the aggregation process, which can be described as $M\leftarrow\sum^{N}_{n=1}w^{n}M^{n}$. These four steps are iteratively executed until the model converges or a predefined number of FL communication rounds $R$ is reached.
@@ III Related Work
IV Proposed Method
IV-A LW-FedMML
The main idea of layer-wise training is to break down the training process of a model into multiple stages, denoted as $s\in[1,S]$. Initially, a given encoder $F$ starts off without any layers, essentially as an empty framework. At the beginning of each stage $s$, a new layer $L^{s}$ is sequentially attached to the encoder $F$, thereby increasing its depth. During stage $s$, only the corresponding layer $L^{s}$ is trained, while the other layers $[L^{1},\dots,L^{(s-1)}]$ from previous stages are frozen without receiving updates. Each stage $s$ lasts for a specified number of epochs, or communication rounds in the context of FL.
Fig. 3 illustrates the training process for LW-FedMML. Specifically, given different modalities $a$ and $b$, layers $L^{s}_{a}$ and $L^{s}_{b}$ are attached to encoders $F_{a}$ and $F_{b}$ at the beginning of each stage $s$. For LW-FedMML without labels, we follow the CLIP [3] approach, which computes the InfoNCE [34] loss for each modality. This approach allows for effective alignment of representations from different data modalities with simplicity. Give a batch of samples $x_{a}$ and $x_{b}$, the projections are extracted as $z_{a}=H_{a}(F_{a}(x_{a}))$ and $z_{b}=H_{b}(F_{b}(x_{b}))$, where $H_{a}$ and $H_{b}$ denote modality-specific projection heads. When both $z_{a}$ and $z_{b}$ are normalized, the loss for modality $a$ can be calculated as follows:
where $B$ is the batch size and $\tau$ represents the temperature parameter. This loss is designed to promote the cosine similarity between representations of different modalities originating from the same sample. Simultaneously, it pushes the representations from different samples away from each other. Following the same principles, the loss for modality $b$ can be calculated as $\ell(z_{b},z_{a})$. The final training loss can be defined as:
$\ell_{clip}(z_{a},z_{b})=\frac{1}{2}(\ell(z_{a},z_{b})+\ell(z_{b},z_{a})).$ (4)
In supervised LW-FedMML, the concatenated representations of $z_{a}$ and $z_{b}$ are passed through a prediction head $H_{\text{sup}}$ for making predictions, as shown in Fig 4. The cross-entropy loss is calculated between the prediction and the ground truth label for training.
@@ IV-A LW-FedMML
2:$L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow$ Download $L_{a}^{s},L_{b}^{s},H_{a},H_{b}$
3:$F_{a}^{s,n}\leftarrow$ Freeze $[L_{a}^{1},\dots,L_{a}^{(s-1)}]$ and attach $L_{a}^{s,n}$
4:$F_{b}^{s,n}\leftarrow$ Freeze $[L_{b}^{1},\dots,L_{b}^{(s-1)}]$ and attach $L_{b}^{s,n}$
5:for epoch $e=1,2,\dots,E$ do
6:     for each batch $(x_{a},x_{b})\in D^{n}$ do
7:         $z_{a}\leftarrow H_{a}^{n}(F_{a}^{s,n}(x_{a}))$
8:         $z_{b}\leftarrow H_{b}^{n}(F_{b}^{s,n}(x_{b}))$
9:         $\mathcal{L}\leftarrow\ell_{\text{clip}}(z_{a},z_{b})$
10:         $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow$ Update with $\nabla\mathcal{L}$
11:     end for
@@
13:Upload $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}$
TABLE I: Comparison of layer management strategies across different training approaches.
@@
TABLE II: Distribution of FL communication rounds for each stage of layer-wise training on the COCO dataset.
Rounds $\boldsymbol{R_{s}}$ for Stage $\boldsymbol{s}$ Total Number of Rounds Text-to-Image Image-to-Text
@@
TABLE III: Learning rate for encoders of different modalities.
@@
FL clients perform the aforementioned steps as the local training process. The detailed procedure of LW-FedMML is shown in Algorithms 1 and 2. Algorithm 1 describes the server-side execution, responsible for determining the stage $s$ and initializing new layers $L_{a}^{s}$ and $L_{b}^{s}$ for each modality, as well as completing the aggregation process. On the other hand, the clients execute Algorithm 2, where they train the active layers $L_{a}^{s}$ and $L_{b}^{s}$ along with the projection heads $H_{a}$ and $H_{b}$. For the supervised setting, clients also train the prediction head $H_{\text{sup}}$ and exchange it with the server. By training only the active layers in each stage, memory and computational resource requirements for clients are significantly reduced. This enables resource-limited client devices to actively participate in the FL process, allowing to take advantage of their private data for model training. Furthermore, only the active layers are required to exchange between the clients and the server. Consequently, both upload and download costs for model exchange are reduced, minimizing FL communication bottlenecks.
IV-B Prog-FedMML
Inspired by [31], we also explore the progressive training approach for federated multimodal learning, denoted as Prog-FedMML. Progressive training, as demonstrated in [31] and [32], has the potential to outperform end-to-end training. Similar to the layer-wise approach, progressive training begins with an initial encoder $F$ without any layers. A new layer $L^{s}$ is sequentially added at each stage $s$. However, unlike the layer-wise approach, all existing layers in $F$ are trained during each stage $s$. Specifically for multimodal learning, Prog-FedMML trains layers $[L_{a}^{1},\dots,L_{a}^{s}]$, and $[L_{b}^{1},\dots,L_{b}^{s}]$ within encoders $F_{a}$ and $F_{b}$ during stage $s$. As a result, Prog-FedMML is more resource-intensive than LW-FedMML, as it requires training all layers added in prior stages, rather than freezing them. The computational and communication resource requirements of Prog-FedMML for an FL training round gradually increase and eventually match those of end-to-end training by the final stage (i.e., $s=S$). This similarity in resource demands can lead to the same drawbacks as end-to-end training, particularly in excluding clients with limited resources from the FL process, thereby losing valuable data for model training. We compare the different characteristics of conventional end-to-end, progressive, and layer-wise training approaches for an encoder $F$ in Table I, highlighting how the layers are managed and utilized at each stage.

### v2

@@ Abstract
Combining different data modalities enables deep neural networks to tackle complex tasks more effectively, making multimodal learning increasingly popular. To harness multimodal data closer to end users, it is essential to integrate multimodal learning with privacy-preserving approaches like federated learning (FL). However, compared to conventional unimodal learning, multimodal setting requires dedicated encoders for each modality, resulting in larger and more complex models. Training these models requires significant resources, presenting a substantial challenge for FL clients operating with limited computation and communication resources. To address these challenges, we introduce LW-FedMML, a layer-wise federated multimodal learning approach which decomposes the training process into multiple stages. Each stage focuses on training only a portion of the model, thereby significantly reducing the memory and computational requirements. Moreover, FL clients only need to exchange the trained model portion with the central server, lowering the resulting communication cost. We conduct extensive experiments across various FL and multimodal learning settings to validate the effectiveness of our proposed method. The results demonstrate that LW-FedMML can compete with conventional end-to-end federated multimodal learning (FedMML) while significantly reducing the resource burden on FL clients. Specifically, LW-FedMML reduces memory usage by up to 2.7×2.7$\times$2.7 ×, computational operations (FLOPs) by 2.4×2.4$\times$2.4 ×, and total communication cost by 2.3×2.3$\times$2.3 ×. We also explore a progressive training approach called Prog-FedMML. While it offers lesser resource efficiency than LW-FedMML, Prog-FedMML has the potential to surpass the performance of end-to-end FedMML, making it a viable option for scenarios with fewer resource constraints.
@@ I Introduction
Deep neural networks have become an integral part of many practical applications in our daily lives. While these models have traditionally been trained on a single modality, recent advancements in the field are driving a shift towards multimodal learning [1, 2, 3]. This transition is motivated by the realization that leveraging multiple types of data simultaneously can significantly enhance the performance of the models [4]. Just as animal intelligence relies on multiple senses to perceive their environment, combining different data modalities allows models to tackle complex tasks more effectively. Integrating diverse data sources, such as vision and audio, ensures that more information is available to the model, thereby improving its perception and decision-making process [5].
Multimodal learning based on instance discrimination approaches aligns representations of different modalities within the same data sample, thereby learning transferable representations [3, 6]. On the other hand, supervised multimodal learning aims to fuse information from various modalities to enhance the predictive performance of the model [7, 8, 9]. As mobile and IoT devices become increasingly powerful, they are integrated with a broader range of sensors [10]. For instance, smartphones are equipped with cameras, gyroscopes, accelerometers, and more, while smart home IoT devices can include an array of modular sensors for various functionalities. These edge devices capture and store real-world multimodal data, which is valuable in training AI models for practical applications. However, conventional multimodal learning methods require collecting this data into a central repository for training, raising data privacy concerns.
To harness the distributed data while also considering privacy issues, numerous studies have explored decentralized training approaches like federated learning (FL) [11, 12, 13, 14]. In an FL system, the central server collects model parameters trained on the private data of edge devices rather than the data itself [11]. Such a mechanism effectively guarantees the confidentiality of data held by different parties involved in the training process. However, this also means that the FL training process relies heavily on the computation and communication capabilities of FL clients, as they are responsible for model training [15].
@@
During training, reducing the batch size may appear to be a straightforward solution to lower the computational demands on resource-constrained devices. However, in instance discrimination-based multimodal learning approaches (e.g., CLIP [3]), a small batch size may compromise the effectiveness of the contrastive loss, degrading the quality of learned representations [17, 18]. These methods benefit from a larger batch size to increase the number of negative samples, which is essential for the contrastive loss to perform well. Furthermore, simply reducing the batch size only addresses the computational concerns but fails to alleviate the communication burden on FL clients, as they still need to exchange the same model size with the server.
@@
Our contributions are summarized as follows: (i) We introduce LW-FedMML, a layer-wise federated multimodal learning strategy designed to address the resource limitations of FL clients engaged in multimodal tasks. LW-FedMML can match the performance of conventional end-to-end training while significantly reducing resource requirements. (ii) We also explore a progressive training approach, referred to as Prog-FedMML. Although Prog-FedMML does not offer the same level of resource efficiency as LW-FedMML, it has the potential to outperform conventional end-to-end training while maintaining similar memory usage and lowering computational and communication costs. (iii) To validate the effectiveness of our proposed methods, we conduct extensive experiments across various FL scenarios and multimodal learning setups.
@@ II-A Federated Learning
A federated learning (FL) process typically involves a central server and a set of FL client devices [11, 19, 20]. These devices are often found close to end-users, capturing and holding valuable data for model training. Meanwhile, the central server oversees and monitors the FL procedure. Given $N$ clients, the goal of FL is to train a global model $M$ by solving the following optimization problem:
@@
with $|D^{n}|$ and $|D|$ representing the number of data samples in $D^{n}$ and $D$, respectively. As illustrated in Fig. 1, four key steps comprise a standard FL communication round. (i) The central server distributes the global model $M$ to each participating client $n\in[1,N]$. (ii) Each client $n$ trains the received model on its local dataset $D^{n}$ to produce the local model $M^{n}$. (iii) The trained local models $[M^{1},\dots,M^{N}]$ are sent back to the server. (iv) The server aggregates the received local models and updates the global model. The FedAvg [11] algorithm is commonly used for the aggregation process, which can be described as $M\leftarrow\sum^{N}_{n=1}w^{n}M^{n}$. These four steps are iteratively executed until the model converges or a predefined number of FL communication rounds $R$ is reached.
Figure 1: Overview of the federated learning process. (i) The global model $M$ is distributed to clients. (ii) Each client $n$ performs training on the local dataset $D^{n}$. (iii) The local model $M^{n}$ is sent back to the server. (iv) The server performs the aggregation.
@@ II-B Multimodal Learning
(a) Instance discrimination-based multimodal learning.
(b) Supervised multimodal learning.
Figure 2: Common multimodal learning approaches given modalities $a$ and $b$. $M$ denotes the model, and $x$ denotes the input sample. (a) The instance discrimination-based approach focuses on aligning the representations of diverse modalities within a given data sample. Negative samples are omitted in the figure for clarity. (b) The supervised approach aims to fuse information from various modalities, enabling the model to make more informed predictions. Here, $H_{\text{sup}}$ can be any task-specific prediction head.
@@ III Related Work
IV Proposed Method
Figure 3: Overview of LW-FedMML for supervised setting at stage $s$.
IV-A LW-FedMML
The main idea of layer-wise training is to break down the training process of a model into multiple stages, denoted as ${s\in[1,S]}$. Initially, a given encoder $F^{0}$ starts off without any layers. At the beginning of each stage $s$, a new layer $L^{s}$ is sequentially attached to the encoder $F^{s}$, thereby increasing its depth. During stage $s$, only the corresponding layer $L^{s}$ is trained, while the other layers $[L^{1},\dots,L^{(s-1)}]$ from previous stages are frozen without receiving updates. Each stage $s$ lasts for a specified number of epochs, or communication rounds in the context of FL.
Figure 4: Overview of LW-FedMML for the instance discrimination-based setting. The training process is divided into multiple stages $s\in[1,S]$. In stage $s$, the active layers $L_{a}^{s}$ and $L_{b}^{s}$ are depicted in green within encoders $F_{a}^{s}$ and $F_{b}^{s}$, where $a$ and $b$ represent different modalities. Prior layers within the encoders are frozen, as indicated in gray.
@@ IV-A LW-FedMML
2:$L_{a}^{s,n}\leftarrow L_{a}^{s},\quad H_{a}^{n}\leftarrow H_{a}$
3:$L_{b}^{s,n}\leftarrow L_{b}^{s},\quad H_{b}^{n}\leftarrow H_{b}$
4:$F_{a}^{s,n}\leftarrow$ Freeze $[L_{a}^{1},\dots,L_{a}^{(s-1)}]$ and attach $L_{a}^{s,n}$
5:$F_{b}^{s,n}\leftarrow$ Freeze $[L_{b}^{1},\dots,L_{b}^{(s-1)}]$ and attach $L_{b}^{s,n}$
6:for epoch $e=1,2,\dots,E$ do
7:     for each batch $(x_{a},x_{b})\in D^{n}$ do
8:         $z_{a}\leftarrow H_{a}^{n}(F_{a}^{s,n}(x_{a}))$
9:         $z_{b}\leftarrow H_{b}^{n}(F_{b}^{s,n}(x_{b}))$
10:         $\mathcal{L}\leftarrow\ell_{\text{clip}}(z_{a},z_{b})$
11:         $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}\leftarrow$ Update with $\nabla\mathcal{L}$
@@
13:end for
14:Upload $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}$
Fig. 4 illustrates the training process for LW-FedMML. Specifically, given different modalities $a$ and $b$, layers $L^{s}_{a}$ and $L^{s}_{b}$ are attached to encoders $F_{a}^{s}$ and $F_{b}^{s}$ at the beginning of each stage $s$. In instance discrimination-based LW-FedMML, we follow the CLIP [3] approach, which computes the InfoNCE [34] loss for each modality. This approach allows effective alignment of representations from different data modalities with simplicity. Given a batch of paired samples $x_{a}$ and $x_{b}$, the projections are extracted as $z_{a}=H_{a}(F_{a}^{s}(x_{a}))$ and $z_{b}=H_{b}(F_{b}^{s}(x_{b}))$, where $H_{a}$ and $H_{b}$ denote modality-specific projection heads. When both $z_{a}$ and $z_{b}$ are normalized, the loss for modality $a$ can be calculated as follows:
where $B$ is the batch size and $\tau$ represents the temperature parameter. This loss is designed to promote the cosine similarity between representations of different modalities originating from the same sample. Simultaneously, it pushes the representations of different samples away from each other. Following the same principles, the loss for modality $b$ can be calculated as $\ell(z_{b},z_{a})$. The final training loss can be defined as:
$\ell_{clip}(z_{a},z_{b})=\frac{1}{2}(\ell(z_{a},z_{b})+\ell(z_{b},z_{a})).$ (4)
In supervised LW-FedMML, the fused representations of $z_{a}$ and $z_{b}$ are passed through a prediction head $H_{\text{sup}}$ for making predictions, as shown in Fig 3. The cross-entropy loss is calculated between the prediction $\hat{y}$ and the ground truth label $y$ for training.
TABLE I: Comparison on how the layers are managed across different training approaches.
@@
FL clients perform the aforementioned steps as the local training process. The detailed procedure of LW-FedMML is shown in Algorithms 1 and 2. Algorithm 1 describes the server-side execution, responsible for determining the stage $s$ and initializing new layers $L_{a}^{s}$ and $L_{b}^{s}$ for each modality, as well as completing the aggregation process. On the other hand, the clients execute Algorithm 2, where they train the active layers $L_{a}^{s}$ and $L_{b}^{s}$ along with the projection heads $H_{a}$ and $H_{b}$. For the supervised setting, clients also train the prediction head $H_{\text{sup}}$ and exchange it with the server. In Appendix A, we describe the detailed procedure for supervised LW-FedMML using Algorithms 3 and 4. By training only the active layers in each stage, memory and computational resource requirements for clients are significantly reduced. This enables resource-limited client devices to actively participate in the FL process, allowing to take advantage of their private data for model training. Furthermore, only the active layers are required to exchange between the clients and the server. Consequently, both upload and download costs for model exchange are reduced, minimizing FL communication bottlenecks.
IV-B Prog-FedMML
Inspired by [31], we also explore the progressive training approach for federated multimodal learning, denoted as Prog-FedMML. Similar to the layer-wise approach, progressive training begins with an initial encoder $F^{0}$ without any layers. A new layer $L^{s}$ is sequentially added at each stage $s$. However, unlike the layer-wise approach, all existing layers in $F^{s}$ are trained during each stage $s$. Specifically for multimodal learning, Prog-FedMML trains layers $[L_{a}^{1},\dots,L_{a}^{s}]$ and $[L_{b}^{1},\dots,L_{b}^{s}]$ within encoders $F_{a}^{s}$ and $F_{b}^{s}$ during stage $s$. Progressive training, as demonstrated in [31] and [32], has the potential to outperform end-to-end training. This advantage likely stems from training and increasing the model depth step-by-step, which may reduce the risk of vanishing or exploding gradients and allow earlier layers to receive more effective gradient updates. This phenomenon is examined further in Section V-C.
Prog-FedMML is more resource-intensive compared to LW-FedMML, as it requires training all layers added in prior stages rather than freezing them. The computational and communication resource requirements of Prog-FedMML for an FL training round gradually increase and eventually match those of end-to-end training by the final stage (i.e., when $s=S$). This similarity in resource demands could lead to the same drawbacks as end-to-end training, particularly in excluding clients with limited resources from the FL process, thereby losing valuable data for model training. We compare the different characteristics of end-to-end, progressive, and layer-wise training approaches for an encoder $F$ in Table I, highlighting how the layers are managed and utilized at each stage.
V Experiment
V-A Experimental Setup
Dataset. We conduct our experiments for multimodal learning in two settings: an instance discrimination-based setting and a supervised setting with labeled data. We use the COCO [35] dataset for the instance discrimination-based setting. This dataset contains 118,287 images for training and 5,000 images for testing, with each image paired with at least five captions. We evaluate the performance on text-to-image and image-to-text retrieval tasks. For the supervised setting, we use the ADVANCE [9] dataset for audiovisual aerial scene recognition. This dataset consists of 5,075 aerial images accompanied by their respective audio recordings, categorized into 13 different scene classes. The underlying premise is that geographic locations are associated with certain sound events, and utilizing both aerial images and sound can lead to more accurate predictions [9]. We allocate 20% of the ADVANCE dataset for testing.
TABLE II: Distribution of FL communication rounds to each stage of layer-wise training on the COCO dataset.
Rounds $\boldsymbol{R^{s}}$ for Stage $\boldsymbol{s}$ Total Number of Rounds Text-to-Image Image-to-Text
@@ V-A Experimental Setup
TABLE III: Learning rates for encoders of different modalities.
@@
We uniformly split the data among the clients for the COCO dataset. We set the batch size to 256 and use $\tau=0.05$ in Eq. 3. For LW-FedMML and Prog-FedMML, we use the learning rates listed in Table III for the image and text encoders at each stage. These values are guided by the transfer learning principles, where pre-trained early layers capture generic features and require minimal fine-tuning, while later layers become more task-specific. As a result, we use lower learning rates for the early stages and gradually increase them for the later stages. The learning rate for the projection heads of both modalities is set to 1e-4. For the conventional FedMML, the learning rates for the image and text encoders are set to 1e-7 and 1e-5, respectively, while the projection heads have a learning rate of 1e-4. Unless otherwise stated, we use the COCO dataset for our experiments.
[…]
V-B Instance Discrimination-based Setting
In Table IV, we compare the performance (Recall@K) of different approaches and the relative resource requirements for a single client. We measure the maximum memory required during training, while FLOPs and communication costs are measured in total across all local training epochs and FL rounds. For FLOPs calculation, we follow the same strategy as in [32], accounting for only one input sample rather than the entire batch. As the number of samples increases, the FLOPs would also increase drastically.
Table IV demonstrates that LW-FedMML achieves comparable performance to end-to-end FedMML, with a minor drop of $\approx 1\%$ in average recall for text-to-image and image-to-text retrieval tasks, while significantly reducing resource requirements. Specifically, LW-FedMML shows a 2.7×2.7$\times$2.7 × reduction in memory usage, a 2.4×2.4$\times$2.4 × decrease in FLOPs requirements, and 2.3×2.3$\times$2.3 × lower communication costs.
Prog-FedMML achieves the best performance, surpassing end-to-end FedMML, which aligns with findings from prior studies [31, 32]. However, the resource requirements of Prog-FedMML are significantly higher than those of LW-FedMML, closely matching those of end-to-end FedMML in terms of memory and communication costs. This could present similar challenges for Prog-FedMML as those faced by end-to-end FedMML, where many clients may fail to meet the resource requirements for participation, thereby preventing their data from being used in model training. Fig. 6 plots the memory usage, FLOPs, and communication costs during the training process for a client. Across all metrics, LW-FedMML consistently shows significantly lower resource requirements compared to FedMML. In contrast, resource demands for Prog-FedMML gradually increase throughout the training, eventually matching those of FedMML.